Posts

NPP Projection Close 4/5

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Minecraft  played just fine at 6.0 this evening (fortunately the crowds did not seem too rowdy or eager to make noise of their own over the film) but I will continue to keep an ear on levels Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Westward the Women is still on the cart awaiting its return pickup Print/video Inspection: Finished inspecting The Rules of the Game and added more stills to the shared drive Asset tracking doc updated/checked?: Maintenance/cleaning: Regular garbage and recycling were tossed + placed empty/additional bags in the bins in 4/6 Cue sheets updated/checked (every shift): Double checked all cues and cross referenced the LMS schedule, adjusted accordingly, and initialed for confirmation Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by...

NPP Projection Report 4/5 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: With how full these Minecraft screenings are getting, I put a 6.0 volume cue for the feature in all playlists. Could even go higher, depending on the crowd. Deleted all 3/28 playlists Trailer Updates: Special Events/ AV set up: NSF Matinee Two went well! Q&A was recorded and it's in the Google Drive folder here Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested the 5.1 OCAP version of Warfare . I don't think we should play the IAB version, not sure if it would sound bad, it's similar to Atmos. There's no regular 7.1 OCAP version. I deleted the IAB one and put the 5.1 version in the OCAP playlist. Outgoing/Shipping (content): Deleted Eephus , Opus , and Flow from everywhere. Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Swept and mopped the booth 2 floor. Cue sheets updated/checked (every shift): Marquee updated?: Pos...

NPP Projection Close 4/4

Image
To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Return of the Living Dead  keys were delivered and uploaded to the LMS Outgoing/Shipping (content): Print/video Inspection: Continued inspecting The Rules of the Game,  took some still photos and added these to the shared drive, and placed the reels heads out back on their shipping reels (4-6 remain and can be looked at tomorrow evening) Asset tracking doc updated/checked?: Maintenance/cleaning: Sorted through some of the posters behind 3/7 to consolidate stock and make space for future releases + added older and duplicates of posters to the trivia section Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes: Removed outgoing titles from the first run whiteboard in 4/6 Another...

Projection Admin 4/4/25

To do this evening/ was completed: Went through all of the remaining NSF playlists one more time to make sure everything is right. They're all good to go. Looked at 12 Monkeys, all good. Damsels in Distress looked at, all good. Inspected the first reel of Rules of the Game. The inspection report is open on the desktop in T1. Presentation Issues: Preshow Clips/Playlist: Built all new A1 playlists for the week of 4/11, plus the available SE shows:  Gosford Park, Hyenas, Fright Night, Superbad, Sundays on Fire - Lillian Trailer Updates: Trailer packs updated. Ingested the trailer for Friendship  - Lillian Special Events/ AV set up: Set up a 2x1 HDMI switcher and audio extractor in Theater 1 which will allow us to switch between laptop/cable setups without having to move any cables around. It even has a remote, which I wasn't expecting. Another is on the way for Theater 2 and should be here tonight. -A Reached out about upcoming private/special events - Lillian  Ordering/Rece...

NPP Projection Close 4/3

To do this evening/ was completed: Made the rounds throughout the evening and kept on eye and ear on picture and sound (see below) Presentation Issues: Theater 2's bulb is on its way out but, dare I say it, it looks like theater 1 might be suffering from a slight flicker as well? Could use another pair of eyes on it to confirm but this doesn't seem to be affecting the presentation too drastically Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Posters for Downton Abbey: The Grand Finale , Nobody 2 , Bring Her Back , and The Materialists were delivered and stored behind 3/7 Ordering/Receiving (content): Gosford Park feature content was delivered via Deluxe and transferred/ingested to theater 4 for next weekend Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Removed Superhuman Body...  from the...

Projection Admin 4/3

To do this evening/ was completed: Hey everyone! Sent out a slightly belated weekly update, so please get your eyes on it when you can. I've put a reminder in there to read through the instructions I sent yesterday for regaining access to Splashtop on your phone, so be sure to check that out. Thanks. - Lillian School screening in T1 went very well this morning! Mics are back and charging. PLEASE DO NOT remove the feature of Superhuman Body from the LMS as we have one more (potential) screening on 4/28. We just need new KDMs, which we should be receiving later in the month. I'll keep y'all updated. - Lillian Presentation Issues: Preshow Clips/Playlist: Tech'd The Encampments , The Friend , A Minecraft Movie . Adjusted volume and double checked credits info on the cue sheets. - Lillian  Built and scheduled Damsels and Kings of Comedy playlists. -A Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content):  Outg...

NPP Projection Close 4/2

To do this evening/ was completed: NSF Opening Nite went smoothly. Mics returned to the booth. I recorded the Q&A but forgot to transfer it to Google Drive. DJ equipment broken down at the end of the night. I put the mixer in the storage closet and not inside of the turntable case itself just fyi. Jeremy asked to borrow an extension cord from now until Sunday, so he'll return it then. Put up the letters for Superhuman. Center row with "1030 AM" directly beneath it. I couldn't find any marquee letters with a colon printed on them so beware we are technically missing that piece of the title. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for ...